# Towards Automated Identification of Violation Symptoms of Architecture Erosion

### v4

@@ Abstract.
Architecture erosion has a detrimental effect on maintenance and evolution, as the implementation deviates from the intended architecture. To prevent this, development teams need to understand early enough the symptoms of erosion, and particularly violations of the intended architecture. One way to achieve this, is through the automated identification of architecture violations from textual artifacts, and particularly code reviews. In this paper, we developed 15 machine learning-based and 4 deep learning-based classifiers with three pre-trained word embeddings to identify violation symptoms of architecture erosion from developer discussions in code reviews. Specifically, we looked at code review comments from four large open-source projects from the OpenStack (Nova and Neutron) and Qt (Qt Base and Qt Creator) communities. We then conducted a survey and semi-structured interviews to acquire feedback from the involved participants who discussed architecture violations in code reviews, to validate the usefulness of our trained classifiers. The results show that the SVM classifier based on word2vec pre-trained word embedding performs the best with an F1-score of 0.779. In most cases, classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Furthermore, 200-dimensional pre-trained word embedding models outperform classifiers that use 100- and 300-dimensional models. In addition, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperforms the individual classifiers. Finally, the findings derived from the online survey and interviews conducted with the involved developers reveal that the violation symptoms identified by our approaches have practical value and can provide early warnings for impending architecture erosion.
Architecture Violation, Architecture Erosion Symptom, Natural Language Processing, Machine Learning, Deep Learning
@@ 1. Introduction
Software architecture is regarded as critical in developing large and complex systems, as it reflects the system structure and behavior and acts as a bridge between business goals and the implemented system (Bass2021sap, ). However, the phenomenon of architecture erosion has become one of the major challenges of architecting, impeding software maintenance and undermining software sustainability (Li2022SMS, ; Le2016rad, ; Venters2018ss, ). Architecture erosion refers to the increasing and accumulating divergences between implemented and intended architecture (Li2022SMS, ; PerryWolf1992AEr, ). It is often described by various terms in the literature (Li2022SMS, ) and practice (Li2021uae, ), such as architectural decay, degradation, and deterioration.
@@
As evidenced in previous studies (Li2021uae, ; Li2022SMS, ; Herold2015dvc, ; Fontana2016aer, ; Macia2012anomalies, ), various symptoms or signs can be observed when architecture erosion occurs during a software system’s life cycle. Our recent systematic mapping study (Li2022SMS, ) reported four categories of such architecture erosion symptoms: structural symptoms (e.g., cyclic dependencies), violation symptoms (e.g., layering violation), quality symptoms (e.g., high defect rate), and evolution symptoms (e.g., extensive ripple effects of changes).
@@
In this study, we specifically concentrate on violation symptoms, as those are the most immediate symptoms of architecture erosion and they require the most attention from practitioners (Li2022sae, ; Li2023wvs, ). This symptom type stems from the definition of architecture erosion (Li2022SMS, ), as a violation causes the implemented architecture to diverge from the intended one. Architecturally-relevant violations in software systems include violations of design principles, architecture patterns, decisions, and requirements (Li2022SMS, ). The accumulation of violation symptoms can render the architecture completely untenable (DeSilva2012AErsurvey, ). For brevity, we refer to violation symptoms of architecture erosion as violation symptoms throughout the remainder of this paper.
@@
While a few temporary violation symptoms might be innocuous and might not result in software errors, the accumulation of architecture violations can affect architecture erosion (PerryWolf1992AEr, ; DeSilva2012AErsurvey, ; Mendoza2021avd, ); consequently, erosion negatively affects quality attributes, including maintainability and performance (Li2021uae, ; Mendoza2021avd, ). Thus, it is essential to identify and monitor violation symptoms to facilitate the maintenance of the system architecture and further reveal inconsistencies between the implemented and the intended architecture. This is the first step towards, eventually, repairing or mitigating architecture erosion (DeSilva2012AErsurvey, ; Li2022SMS, ).
@@
Violation symptoms can be identified through source code analysis, but that has certain limitations. As pointed out in previous studies (Sharma2018sss, ; Li2022SMS, ; Li2023wvs, ), a number of factors hinder the effective utilization of static code analysis tools in precisely identifying violation symptoms through source code. For example, such tools only detect a limited set of types of architecture violations, they only support few programming languages, and they do not capture the semantics of violation symptoms as they rely solely on code metrics. Alternatively, violations symptoms can be manually identified by analyzing textual artifacts that contain information related to system architecture and design, such as code comments, code reviews, and issue trackers (Li2022sae, ). In this context, recent work (Li2022sae, ; Li2023wvs, ) has introduced a taxonomy on violation symptoms, that can be used as a checklist to detect such symptoms during code review. However, it could be tedious, time-consuming, and potentially error-prone to manually identify potential architecture violations (Li2023wvs, ). Automated techniques could address this problem, but to the best of our knowledge, no prior study has investigated the automated identification of violation symptoms in textual artifacts, such as code review comments.
@@
To address this gap, we conduct an exploratory study to assess the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. The automated identification of violation symptoms could act in a complementary way to modern code review. Modern code review usually addresses code changes in general through knowledge sharing (badampudi2023modern, ), while automated identification of violation symptoms can offer an extra layer of scrutiny. For example, such an automated approach can issue warnings for potential or ignored violation symptoms, preventing the integration of code changes with architectural violations into the code base, and ensuring a more robust and reliable development process. Moreover, automatically identifying violation symptoms from code review comments could offer valuable insights at the architecture level and facilitate the corresponding knowledge sharing between developers and reviewers. The ultimate purpose of our approach is to complement the modern code review process, through warning of and guarding against architectural violation issues, eventually maintaining architectural integrity.
@@
In this work, we developed 15 classifiers based on Machine Learning (ML) and 4 classifiers based on Deep Learning (DL), to identify violation symptoms from developer discussions in code reviews of four large open-source software projects from the OpenStack and Qt communities. Our results show that the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803. We also found out that utilizing fastText pre-trained word embedding can achieve relatively good results and it helps to improve the classifiers’ performance. Furthermore, 200-dimensional pre-trained word embedding models outperform classifiers that use 100- and 300-dimensional models. Finally, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperform the individual classifiers.
We then conducted an online survey and semi-structured interviews to validate the effectiveness of automated identification of violation symptoms in practice. We received responses from 24 survey participants and 6 interviewees, all of whom were involved in the discussion of the collected review comments concerning architecture violations. The results of our survey and interview show that the practitioners’ perception of the classifiers’ results confirms the promising applications of automated identification of violation symptoms in practice. Especially, such classifiers can enable and inspire the participants to find, prioritize, and handle architecturally-related violation issues.
@@
Unlike previous studies that used source code analysis to identify architectural issues (see Section 2.1), our work takes a different approach by focusing on identifying violation symptoms through textual artifacts, which better aligns with how practitioners view these symptoms. In addition, our works explores the possibility of identifying violation symptoms of architecture erosion by using ensemble learning techniques, offering advantages over single ML/DL algorithms. The main contributions of this work are the following:
  • This is the first work to explore the application of ML and DL-based classifiers regarding the automated identification of violation symptoms in code review comments.
  • We compared the performance of several trained classifiers using popular ML and DL-based algorithms, as well as the impact of three pre-trained word embeddings on classification performance.
@@
  • We gained insights from participants regarding the usefulness of ML and DL-based classifiers in identifying architecture violations during code review. We caution against the indiscriminate use of DL techniques, particularly for tasks where return on investment is a primary concern.
@@ 2. Background
In this section, we introduce a) the background of architecture erosion and its symptoms; b) architecture conformance checking, which is the most commonly used approach to detect architecture violations; and c) the modern code review process in Gerrit [^1] https://www.gerritcodereview.com/, as this is used in our dataset.
@@ 2.1. Architecture Erosion and Related Symptoms
In the past decades, a wide variety of studies are concerned with the architecture erosion phenomenon, which has been extensively discussed and described with various terms (Li2021uae, ; Li2022SMS, ), such as architecture decay (Hassaine2012ADvISE, ; Le2018emad, ), degradation (Lenhard2019ess, ; Herold2020asd, ), degeneration (Hochstein2005cad, ).
@@
Table 1. Examples of violation symptoms of architecture erosion collected from Nova, Neutron, Qt Base, and Qt Creator
@@
Qt Base This one seems to be a layering violation: why does QNetworkRequest understand QHttpNetworkConnection?”
This breaks the abstraction layer of publicQtConfig and is annoyingly verbose. instead, publicQtConfig should have an optional parameter ’name’ to override the built-in ‘feature’ member.
@@
As mentioned in the Introduction section, we focus on the most direct erosion symptoms of architecture erosion (i.e., violation symptoms). Table 1 presents several examples of violation symptoms of architecture erosion from the used dataset (see Section 3); a detailed comparison of our work and related work is given in Section 7.
@@ 2.2. Architecture Conformance Checking
In addition, rule-based conformance checking approaches are also employed to identify architecture violations. For example, previous studies identified architecture violations by checking explicitly defined architectural rules (Mendoza2021avd, ; Schroder2017acc, ; Terra2009dcl, ). Moreover, it is viable to check architecture conformance and identify architecture violations by defining and describing the systems through Architecture Description Languages (ADLs) (Terra2009dcl, ; Rocha2017DCL, ), or Domain-Specific Languages (DSLs) (Caracciolo2015uaa, ; Juarez2017pee, ).
Despite their potential, the previously mentioned approaches for identifying architecture violations have clear limitations. For instance, much effort is required to address the challenges of understanding the architecture design (e.g., concepts and relations), defining architectural rules (or description languages) in advance, and establishing a mapping between architectural elements and source code (Li2022SMS, ; Li2023wvs, ). Additionally, other limitations, such as lack of generalizability, lack of visualization of architectural views, and insufficient tooling support, have impeded the wide adoption of these approaches in practice (Li2022SMS, ).
@@ 2.3. Code Review in Gerrit
Code review is a crucial software development activity that involves the systematic examination of assigned code to identify defects and improve software quality (Bacchelli2013eoc, ). A methodical code review process not only enhances the quality of software systems, but also facilitates the sharing of development knowledge and prevents the release of unstable and defective products. Over time, code review practices have become increasingly important, have been widely adopted in modern software development, and have received extensive tool support. In fact, tool-based code review has become the norm in both industry and open-source communities, with a variety of available code review tools, such as Me’s Phabricator [^2] https://www.phacility.com/, VMware’s Review-Board [^3] https://www.reviewboard.org/, and Gerrit.
@@
The two communities (i.e., OpenStack and Qt), from which the four open-source software (OSS) projects were selected in this work, use Gerrit as the code review tool. Figure 1 illustrates the modern code review process in Gerrit, which is a well-known code review tool and widely used in modern code review (Davila2021SLRmcr, ). Gerrit allows for efficient collaboration on code changes before they are merged into the main code base. Once a developer commits new code changes (e.g., software patches) and their descriptions to Gerrit, the system creates a page to record all changes and relevant descriptions (e.g., commit messages). Then, the system conducts a sanity check to verify whether the patch is compliant and to ensure that the code does not contain obvious compilation errors. After the submitted patch passes the sanity check, code reviewers manually examine the patch and provide comments to correct any potential errors, followed by giving a voting score. The comments are subsequently used by the developers to improve the patch and submit it for another review. Finally, the submitted patch is merged into the code repository after passing integration tests, which confirms that there are no issues or conflicts. Alternatively, the patch can be abandoned if the voting is negative.
The above process is used, among others, in revisions that enable the analysis of potentially significant architectural changes. Code review comments provide software developers a finer granularity for investigating architectural changes and violations in their daily development routines (Li2021mpl, ). Meanwhile, developers provide reasoning and rationale for the changes they make in code review comments, which enables code review data to be a valuable source of knowledge for explaining changes (Paixao2021crac, ), offering insights into developers’ concerns, suggestions, and potential violations.

### v5

@@ Abstract.
Architecture erosion has a detrimental effect on maintenance and evolution, as the implementation deviates from the intended architecture. To prevent this, development teams need to understand early enough the symptoms of erosion, and particularly violations of the intended architecture. One feasible way is through the automated identification of architecture violations from textual artifacts, and particularly code reviews. In this paper, we developed 15 machine learning-based and 4 deep learning-based classifiers with three pre-trained word embeddings to identify violation symptoms of architecture erosion from developer discussions in code reviews. Specifically, we looked at code review comments from four large open-source projects from the OpenStack (Nova and Neutron) and Qt (Qt Base and Qt Creator) communities. We then conducted a survey and semi-structured interviews to acquire feedback from the involved participants who discussed architecture violations in code reviews, to validate the usefulness of our trained classifiers. Moreover, we conducted additional comparative experiments by evaluating Large Language Model (LLM)-based classifiers, including GPT-4o, Qwen-2.5, and DeepSeek-R1. The results show that, for ML/DL-based classifiers, the SVM classifier based on word2vec pre-trained word embedding performs the best with an F1-score of 0.779. In most cases, classifiers with the fastText pre-trained word embedding model can achieve relatively good performance. Moreover, 200-dimensional pre-trained word embedding models outperform classifiers that use 100- and 300-dimensional models. For ML/DL-based classifiers, ensemble classifiers based on the majority voting strategy can enhance the classifier and outperform the individual classifiers. The findings derived from the online survey and interviews conducted with the involved developers reveal that the violation symptoms identified by our approaches have practical value and can provide early warnings for impending architecture erosion. Furthermore, LLM-based classifiers consistently outperform traditional ML/DL models, with GPT-4o yielding the highest F1-score of 0.851, though ensemble strategies offered no further performance gains for ensemble LLM-based classifiers in our case. We investigate the automated identification of violation symptoms from code reviews using both traditional ML/DL and state-of-the-art LLM techniques. Our contributions include the automated approach for identifying violation symptoms in code reviews, a systematic comparison of ML/DL and LLM approaches, and practitioner‐centered insights into their practical usefulness, ultimately contributing to better architectural conformance and sustainability in software systems.
Architecture Violation, Architecture Erosion Symptom, Natural Language Processing, Machine Learning, Deep Learning, Large Language Models
We would like to thank all the practitioners who participated in our survey and interviews. This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62402348 and 62172311.
@@ 1. Introduction
Software architecture is critical in developing large and complex systems, as it reflects the system structure and behavior and acts as a bridge between business goals and the implemented system (Bass2021sap, ). However, the implemented architectures often diverge from their intended architectures over time, which is the phenomenon of architecture erosion (Li2022SMS, ; PerryWolf1992AEr, ). This phenomenon undermines maintainability, performance, and overall sustainability (Li2022SMS, ; Le2016rad, ; Venters2018ss, ) and often described by various terms in the literature (Li2022SMS, ) and practice (Li2021uae, ), such as architectural decay, degradation, and deterioration. Various symptoms or signs can be observed when architecture erosion occurs during a software system’s life cycle (Li2021uae, ; Li2022SMS, ; Herold2015dvc, ; Fontana2016aer, ; Macia2012anomalies, ). Our recent systematic mapping study (Li2022SMS, ) reported four categories of such architecture erosion symptoms: structural symptoms (e.g., cyclic dependencies), violation symptoms (e.g., layering violation), quality symptoms (e.g., high defect rate), and evolution symptoms (e.g., extensive ripple effects of changes).
@@
In this study, we specifically focus on violation symptoms, as those are the most immediate symptoms of architecture erosion and require the most attention from practitioners (Li2022sae, ; Li2023wvs, ). This symptom type stems from the definition of architecture erosion (Li2022SMS, ), as a violation causes the implemented architecture to diverge from the intended one. Architecturally-relevant violations in software systems include violations of design principles, architecture patterns, decisions, and requirements (Li2022SMS, ). The accumulation of violation symptoms can render the architecture completely untenable (DeSilva2012AErsurvey, ). For brevity, we refer to violation symptoms of architecture erosion as violation symptoms throughout the remainder of this paper.
@@
While a few temporary violation symptoms might be innocuous and might not result in software errors, the accumulation of architecture violations result in architecture erosion (PerryWolf1992AEr, ; DeSilva2012AErsurvey, ; Mendoza2021avd, ); consequently, erosion negatively affects quality attributes, including maintainability and performance (Li2021uae, ; Mendoza2021avd, ). Thus, it is essential to identify and monitor violation symptoms to facilitate the maintenance of the system architecture and further reveal architecture inconsistencies. This is the first step towards, eventually, repairing or mitigating architecture erosion (DeSilva2012AErsurvey, ; Li2022SMS, ).
@@
Violation symptoms can be identified static analysis of source code, but that has certain limitations. As mentioned in previous studies (Sharma2018sss, ; Li2022SMS, ; Li2023wvs, ), a number of factors hinder the effective utilization of static code analysis tools in precisely identifying violation symptoms through source code. For example, such tools only detect limited violation types and narrow programming language support, and cannot capture semantics of violation symptoms by relying solely on code metrics. Alternatively, violations symptoms can be manually identified by analyzing textual artifacts that contain information related to system architecture and design, such as code comments, code reviews, and issue trackers (Li2022sae, ). Recent studies (Li2022sae, ; Li2023wvs, ) have proposed taxonomies on violation symptoms, that can be used as a checklist to detect such symptoms during code review. However, it could be tedious, time-consuming, and potentially error-prone to manually identify potential architecture violations (Li2023wvs, ). Automated techniques could address this problem, but to the best of our knowledge, no prior study has investigated the automated identification of violation symptoms in textual artifacts, such as code review comments.
@@
To address this gap, our goal is to assess the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. The automated identification of violation symptoms could act in a complementary way to modern code review. Modern code review usually addresses code changes in general through knowledge sharing (badampudi2023modern, ), while automated identification of violation symptoms can offer an extra layer of scrutiny. For example, such an automated approach can issue warnings for potential or ignored violation symptoms, preventing the integration of code changes with architectural violations into the code base, and ensuring a more robust and reliable development process. Moreover, automatically identifying violation symptoms from code review comments could offer valuable insights at the architecture level and facilitate the corresponding knowledge sharing between developers and reviewers. The ultimate purpose of our approach is to complement the modern code review process, through warning of and guarding against architectural violation issues, eventually maintaining architectural integrity.
@@
In this work, we first developed 15 classifiers based on Machine Learning (ML) and 4 classifiers based on Deep Learning (DL), to identify violation symptoms from developer discussions in code reviews of four large open-source software projects from the OpenStack and Qt communities. Our results show that the SVM classifier based on word2vec performed the best with a Precision of 0.789, a Recall of 0.828, an F1-score of 0.808, and an Accuracy of 0.803. We also found out that utilizing fastText pre-trained word embedding can achieve relatively good results and it helps to improve the classifiers’ performance. Moreover, 200-dimensional pre-trained word embedding models outperform classifiers that use 100- and 300-dimensional models. Finally, an ensemble classifier based on the majority voting strategy can further enhance the classifier and outperform the individual classifiers. Then, we conducted an online survey and semi-structured interviews to validate the effectiveness of automated identification of violation symptoms in practice. We received responses from 24 survey participants and 6 interviewees, all of whom were involved in the discussion of the collected review comments concerning architecture violations. Their feedback confirmed the practical value of automated identification of violation symptoms, as such classifiers can support practitioners in finding, prioritizing, and handling architecturally-related issues.
Furthermore, the advent of Large Language Models (LLMs) fundamentally reshaped Software Engineering (SE) paradigms (hou2024large, ; li2025ChatGPT, ), existing studies explored the capacities of LLMs on popular SE tasks (e.g., code generation), but no studies investigate the fine-grained task, identifying violation symptoms in code review comments. Therefore, we further developed LLM-based classifiers using the state-of-the-art LLMs (i.e., GPT-4o, Qwen-2.5, and DeepSeek-R1) and compared their performance with the traditional ML/DL techniques on the same classification task. The results show that LLM-based classifiers typically outperform the traditional ML/DL-based classifiers on all metrics, and GPT-4o-based classifier with 2-shot setting comparatively achieves better performance (F1-score 0.851) than Qwen-2.5 and DeepSeek-R1 on identifying violation symptoms. Notably, ensembling LLM-based classifiers does not improve the performance.
@@
Unlike previous studies that relied on source code analysis to identify architectural issues (see Section 2.1), our work takes a complementary approach by identifying violation symptoms from textual artifacts, which better aligns with practitioners’ perspectives. We also explore ensemble learning techniques to detect architecture erosion symptoms, showing advantages over single ML/DL algorithms. The main contributions of this work are the following:
  • We conducted exploratory experiments to demonstrate the feasibility of ML/DL, LLM-based classifiers for automatically identifying violation symptoms in code review comments, and we compared their performance under different settings, respectively.
  • We investigated the impact of pre-trained word embeddings and shot settings on classification performance. For ML/DL-based classifiers, the SVM-based classifier with a 200-dimensional pre-trained word2vec performs best. Besides, LLM-based classifiers indeed outperform ML/DL-based classifiers, and the GPT-4o-based classifier with a 2-shot setting achieves the best performance.
@@
  • We gained insights from the survey and interview participants regarding the usefulness of trained classifiers in identifying architecture violations during code review. We provide the implications and suggestions based on our empirical study.
@@ 2. Background
In this section, we introduce a) the background of architecture erosion and its symptoms; b) architecture conformance checking, the most commonly used approach to detect architecture violations; and c) the modern code review process in Gerrit [^1] https://www.gerritcodereview.com/.
@@ 2.1. Architecture Erosion and Related Symptoms
In the past decades, a wide variety of studies are concerned with the architecture erosion phenomenon, which has been extensively discussed and described with various terms (Li2021uae, ; Li2022SMS, ), such as architecture decay (Hassaine2012ADvISE, ; Le2018emad, ), degradation (Herold2020asd, ), degeneration (Hochstein2005cad, ).
@@
Table 1. Examples of violation symptoms collected from Nova, Neutron, Qt Base, and Qt Creator
@@
Qt Base I think it’s a layer isolation violation to just make the call here. It should be fixed at the Compute API level rather IMHO.
Feels like this DB work violates the level of abstraction we are expecting here.
@@
As mentioned in the Introduction section, we focus on the most direct erosion symptoms of architecture erosion (i.e., violation symptoms). Table 1 presents several examples of violation symptoms of architecture erosion from the used dataset (see Section 3). To facilitate the examination of additional examples, we have made our dataset publicly available online (onlinepackage, ); a detailed comparison of our work and related work is given in Section 7.
@@ 2.2. Architecture Conformance Checking
In addition, rule-based conformance checking approaches are also employed to identify architecture violations. For example, previous studies identified architecture violations by checking explicitly defined architectural rules (Mendoza2021avd, ). Moreover, it is viable to check architecture conformance and identify architecture violations by defining and describing the systems through Architecture Description Languages (ADLs), or Domain-Specific Languages (DSLs).
Despite their potential, the previously mentioned approaches for identifying architecture violations have clear limitations. For instance, much effort is required to address the challenges of understanding the architecture design (e.g., concepts and relations), defining architectural rules (or description languages) in advance, and establishing a mapping between architectural elements and source code (Li2022SMS, ; Li2023wvs, ). Additionally, other limitations, such as a lack of generalizability and architectural views visualization, insufficient tooling support, have impeded the wide adoption of these approaches in practice (Li2022SMS, ).
@@ 2.3. Code Review in Gerrit
Code review is a crucial software development activity that involves the systematic examination of assigned code to identify defects and improve software quality (Bacchelli2013eoc, ). A methodical code review process not only enhances the quality of software systems, but also facilitates the sharing of development knowledge and prevents the release of unstable and defective products. Over time, code review practices have become increasingly important, have been widely adopted in modern software development, and have received extensive tool support. In fact, tool-based code review has become the norm in both industry and open-source communities, with a variety of available code review tools, such as Gerrit, GitHub’s Pull Requests, Me’s Phabricator [^2] https://www.phacility.com/, and VMware’s Review-Board [^3] https://www.reviewboard.org/.
@@
The two communities (i.e., OpenStack and Qt), from which the four open-source software (OSS) projects were selected in this work, use Gerrit as the code review tool. Figure 1 illustrates the modern code review process in Gerrit, which is a well-known code review tool and widely used in modern code review (Davila2021SLRmcr, ). Gerrit allows for efficient collaboration on code changes before they are merged into the main code base. Once a developer commits new code changes (e.g., software patches) and their descriptions to Gerrit, the system creates a page to record all changes and relevant descriptions (e.g., commit messages). Then, the system conducts a sanity check to verify whether the patch is compliant and to ensure that the code does not contain obvious compilation errors. After the submitted patch passes the sanity check, code reviewers manually examine the patch and provide comments to correct any potential errors, followed by giving a voting score. The comments are subsequently used by the developers to improve the patch and submit it for another review. Finally, the submitted patch is merged into the code repository only after passing integration tests (potentially over multiple iterations), which confirms that there are no issues or conflicts. Alternatively, the patch can be abandoned if the voting is negative.
The above process is used, among others, in revisions that enable the analysis of potentially significant architectural changes. Code review comments provide developers a finer granularity for investigating architectural changes and violations in their daily development routines (Li2021mpl, ). Meanwhile, developers provide reasoning and rationale for the changes they make in code review comments, which enables code review data to be a valuable source of knowledge for explaining changes (Paixao2021crac, ), offering insights into developers’ concerns, suggestions, and potential violations.